OpenStudy (anonymous):

simplify: (2x^2y^-5)(6x^-3y)(1/3x^-1y^3)

OpenStudy (kira_yamato):

Please retype the question using the equation "service" provided... Thanks...

OpenStudy (anonymous):

\[\left( 2x ^{2}y ^{-5} \right)\left( 6x ^{-3}y \right)\ \left( 1\div3x ^{-1}y ^{3} \right)\]

OpenStudy (kira_yamato):

Hope this helps...

OpenStudy (anonymous):

thank you

OpenStudy (kira_yamato):

My pleasure...

OpenStudy (anonymous):

can you help me with the others? i have 2 more posted!